A streched string 80 cm long has fundamental frequency 300 vibrations per second, when length of string is changed to 40 cm then what will be fundamental frequency? (a) 600 hz (b) 300 hz (c)150 under root 2 ( d)150

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Ajay
 
fundamental frequency of vibration = f , L = legth of the string = 80 cm = 0.80 m 
T ension in the string = T , m = mass per unit length of the string 
f = (1 / 2L) * √ (T / m) 
T and m remaining same, f propotional to (1/L)
(f 1) / (f 2) = (L2) / (L1) ; f1 = 300 Hz, L1 = 80 cm, L2 = 40 cm, f2 = to be found out 
(f 2) / (f 1) = (L1) / (L2) ==> (f 2) = (L1 / L2) * (f 1) = (80/40)*(300) Hz = 600 Hz
 
 
Regards
Arun (askIITians forum expert)